ice cream scoops are often 1.25 inches in diameter. how many scoops should you get from a half gallon of ice cream? a half gallon container can be approximated by a cylinder with a diameter of 4.8 inches and a height of 6.5 inches.: ice cream scoops are often 1.25 inches in diameter. how many scoops should you get from a half gallon of ice cream? a half gallon container can be approximated by a cylinder with a diameter of 4.8 inches and a height of 6.5 inches.
Answer by nerdybill(1040)

Container:
r = 4.8/2 = 2.4 inches
h = 6.5
volume of cylinder = h(pi)r^2
volume of cylinder = 6.5(pi)(2.4)^2
.
cone:
r = 1/2 d = 1/2(1.25) = 0.625 inches
volume of sphere = 4/3(pi)r^3
volume of sphere = 4/3(pi)(0.625)^3
.
# of scoops = "volume of cylinder"/"volume of sphere"
[6.5(pi)(2.4)^2]/[4/3(pi)(0.625)^3]
[6.5(2.4)^2]/[4/3(0.625)^3]
[37.44]/[0.32552]
115.016
.
or, approximately
115 scoops